# Bandit and Delayed Feedback in Online Structured Prediction

### v1

@@ E.1 Details of Optimistic Delayed Adaptive FTRL (ODAFTRL)
Let $\alpha=\frac{B^{2}}{2}$ and $\delta\in(0,1)$. Then, ODAFTRL with the AdaHedgeD update in online structured prediction with a delay of $D$ achieves
$\mathcal{R}_{T}\leq 4BDL+\frac{\sqrt{2}}{3}\log\frac{1}{\delta}+\frac{\left(\sqrt{(1-a)\log\frac{1}{\delta}}+\sqrt{2bB^{2}}\right)^{2}}{a}+\frac{3}{2}\left(a^{-1}bB^{4}L^{2}(D+1)^{2}T\right)^{1/3},$
@@
Let $Z_{t}=L_{t}(\widehat{\bm{y}}_{t})-\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]$. Then, we have $|Z_{t}|\leq 1$ and $\mathbb{E}_{t}\left[{Z_{t}^{2}}\right]\leq(1-a)S_{t}(\bm{W}_{t})$ from Assumption 4.1. Hence, from Lemma D.1, with probability at least $1-\delta$, the first term in (14) is upper bounded as
@@
From from Assumption 4.1 and Lemma E.1, the second term in (14) is also upper bounded as
@@
$\leq 2\left(2BDL+\sqrt{bB^{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}+\left(bB^{4}L^{2}(D+1)^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}\right)-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})},$ (16)
where we used the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$. Therefore, substituting (15) and (16) into (14) gives
$\mathcal{R}_{T}$ $\leq 4BDL+\frac{\sqrt{2}}{3}\log\frac{1}{\delta}+\left(\sqrt{2(1-a)\log\frac{1}{\delta}}+2\sqrt{bB^{2}}\right)\sqrt{\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}$
$\quad+2\left(bB^{4}(D+1)^{2}L^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}$
$\leq 4BDL+\frac{\sqrt{2}}{3}\log\frac{1}{\delta}+\frac{\left(\sqrt{(1-a)\log\frac{1}{\delta}}+\sqrt{2bB^{2}}\right)^{2}}{a}+\frac{3}{2}\left(a^{-1}bB^{4}L^{2}(D+1)^{2}T\right)^{1/3},$
@@
Appendix F Omitted Details of Section 5
@@ F.1 Common analysis
We provide the analysis that is commonly used in the proofs of Theorems 5.1 and 5.2. We use ODAFTRL with the AdaHedgeD update in Section E.1 for $\mathsf{ALG}$. From Lemma E.1, it holds that
$\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\leq 2\left(2\max_{t\in[T]}a_{t-D:t-1}+\sqrt{\sum_{t=1}^{T}{a_{t}^{2}}+B^{2}b_{t}}\right),$ (17)
@@
$a_{t}=B\min\{\|\widetilde{\bm{G}}_{t-D:t}\|_{\mathrm{F}},\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}\}\quad\mbox{and}\quad b_{t}\leq\min\left\{\frac{1}{2}\|\widetilde{\bm{G}}_{t-D:t}\|_{\mathrm{F}}^{2},\|\widetilde{\bm{G}}_{t-D:t}\|_{\mathrm{F}}\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}\right\}.$
@@
$\mathbb{E}\left[\max_{t\in[T]}a_{t-D:t-1}\right]\leq B\mathbb{E}\left[\max_{t\in[T]}\sum_{s=t-D}^{t-1}\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}\right]\leq B\mathbb{E}\left[\sqrt{D\max_{t\in[T]}\sum_{s=t-D}^{t-1}\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}^{2}}\right]\leq B\sqrt{D\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]},$ (18)
from the Cauchy–Schwarz inequality and Jensen’s inequality. Hence, from (17) and (18), we have
$\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]$ $\leq 2\left(2\mathbb{E}\left[\max_{t\in[T]}a_{t-D:t-1}\right]+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{a_{t}^{2}}\right]}+B\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]}\right)$
$\leq 2B\left(2\sqrt{D\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]}\right),$ (19)
@@
$\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t-D:t}\|_{\mathrm{F}}\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}}\right]\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}\sum_{s=t-D}^{t-1}\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}}\right]$
$=\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\mathbb{E}\left[\sum_{t=1}^{T}{\mathbb{E}_{t}\left[\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}\right]\sum_{s=t-D}^{t-1}\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}}\right],$ (20)
@@
$\mathbb{E}\left[\mathcal{R}_{T}\right]\leq\frac{4bB}{a}\left(D^{1/4}+D^{-1/4}\right)^{2}\sqrt{KT}+2BL\sqrt{DT}+B\sqrt{DKT}=O(D\sqrt{KT}).$
@@
$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+L\mathbb{E}\left[\sum_{t=1}^{T}{\sum_{s=t-D}^{t-1}\mathbb{E}_{s}\left[\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}\right]}\right]$
$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+DL^{2}T,$ (21)
where the second and third inequality follow from the inequality $\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}}\right]=\|\bm{G}_{t}\|_{\mathrm{F}}\leq L$. Hence, from (F.1) and (F.2), it holds that
$\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]$ $\leq 2B\left(2\sqrt{D\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+DL^{2}T}}\right)$
$\leq 4B\left(\sqrt{D}+1\right)\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+2BL\sqrt{DT}$
$\leq 4B\left(\sqrt{D}+1\right)\sqrt{\frac{bK}{q}\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}+2BL\sqrt{DT},$ (22)
@@
$\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]=\frac{\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}{p_{t}(\bm{y}_{t})}\leq\frac{K}{q}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq\frac{bK}{q}S_{t}(\bm{W}_{t}).$
@@
$\leq 4B\left(\sqrt{D}+1\right)\sqrt{\frac{bK}{q}\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}+2BL\sqrt{DT}-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
$\leq\frac{4bB^{2}}{a}\left(\sqrt{D}+1\right)^{2}\frac{bK}{q}+2BL\sqrt{DT}+qT,$
where the first inequality follows from Assumption 3.2, the second inequality follows from the convexity of $S_{t}$ and the unbiasedness of $\widetilde{\bm{G}}_{t}$, and the last inequality follows from $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$. Finally, by substituting $q=B\sqrt{DK/T}$, we obtain
$\mathbb{E}\left[\mathcal{R}_{T}\right]\leq\frac{4bB}{a}\left(D^{1/4}+D^{-1/4}\right)^{2}\sqrt{KT}+2BL\sqrt{DT}+B\sqrt{DKT},$
@@
$\mathbb{E}[\mathcal{R}_{T}]=\frac{8bB^{2}}{a}\Big{(}\sqrt{D}+1\Big{)}^{2}+2B(L+C_{x}C_{y})\sqrt{DT}+O\left(B^{2/3}D^{1/3}T^{2/3}\right).$
@@
$\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}}\right]$ $=\mathbb{E}_{t}\left[{\|(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\widetilde{\bm{y}}_{t})\bm{x}_{t}^{\top}\|_{\mathrm{F}}}\right]\leq\mathbb{E}_{t}\left[{\|\bm{G}_{t}\|_{\mathrm{F}}+C_{x}\|\bm{y}_{t}-\widetilde{\bm{y}}_{t}\|_{2}}\right]\leq\|\bm{G}_{t}\|_{\mathrm{F}}+C_{x}\mathbb{E}_{t}\left[{\|\bm{y}_{t}\|_{2}+\|\widetilde{\bm{y}}_{t}\|_{2}}\right]$
@@
$\leq\|\bm{G}_{t}\|_{\mathrm{F}}+C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q}\leq L+C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q},$
where the first inequality follows from $C_{x}=\mathrm{diam}(\mathcal{X})$, the third inequality follows from $C_{y}=\mathrm{diam}(\mathcal{Y})$, the fourth inequality follows from Jensen’s inequality, and the fifth inequality follows from Lemma D.5. Thus, combining (F.1) with the last inequality, we have
@@
$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\left(L+C_{x}C_{y}+\sqrt{\frac{C_{x}^{2}\omega}{q}}\right)\mathbb{E}\left[\sum_{t=1}^{T}{\sum_{s=t-D}^{t-1}\mathbb{E}_{s}\left[\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}\right]}\right]$
$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+DT\left(L+C_{x}C_{y}+\sqrt{\frac{C_{x}^{2}\omega}{q}}\right)^{2}.$ (23)
Hence, from (F.1) and (23), we have
@@
$\leq 2B\left(2\sqrt{D\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]}\right)$
$\leq 4B\left(\sqrt{D}+1\right)\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+2B\left(L+C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q}\right)\sqrt{DT}$
$\leq 4B\left(\sqrt{D}+1\right)\sqrt{2\sum_{t=1}^{T}{\left(bS_{t}(\bm{W}_{t})+\frac{C_{x}^{2}\omega}{q}\right)}}+2B\left(L+C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q}\right)\sqrt{DT},$ (24)
where the first inequality follows from (F.1), the second inequality follows from (23) and the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$, and the last inequality follows from Lemma D.6. Therefore, combining all the above arguments yields
@@
$\leq 4B\left(\sqrt{D}+1\right)\left(\sqrt{2b\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}+\sqrt{2C_{x}^{2}\omega T/q}\right)$
$\qquad+2B\left(L+C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q}\right)\sqrt{DT}-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
$\leq\frac{8bB^{2}}{a}\Big{(}\sqrt{D}+1\Big{)}^{2}+2B(L+C_{x}C_{y})\sqrt{DT}+2BC_{x}\big{(}(2\sqrt{2}+1)\sqrt{D}+2\sqrt{2}\big{)}\sqrt{\omega T/q}+qT,$

### v2

@@ E.3 Proof of high probability bound
Let $\alpha=\frac{B^{2}}{2}$ and $\delta\in(0,1)$. Then, ODAFTRL with the AdaHedgeD update in online structured prediction with a fixed delay of $D$ achieves
$\mathcal{R}_{T}\leq$ $4BC_{x}C_{y}D+\frac{\sqrt{2}}{3}\log\frac{1}{\delta}$
$+\frac{\left(\sqrt{(1-a)\log\frac{1}{\delta}}+\sqrt{2bB^{2}}\right)^{2}}{a}+\min\left\{\frac{b(D+1)^{2}}{a},\frac{3}{2}\left(a^{-1}bB^{4}C_{x}^{2}C_{y}^{2}(D+1)^{2}T\right)^{1/3}\right\},$
@@
Let $Z_{t}=L_{t}(\widehat{\bm{y}}_{t})-\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]$. Then, we have $|Z_{t}|\leq 1$ and $\mathbb{E}_{t}\left[{Z_{t}^{2}}\right]\leq\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]\leq(1-a)S_{t}(\bm{W}_{t})$ from Assumption 4.1. Hence, from Lemma D.1, with probability at least $1-\delta$, the first term in (11) is upper bounded as
@@
From Assumption 4.1 and Lemma E.1, the second term in (11) is also upper bounded as
@@
$\leq 4BC_{x}C_{y}D+\sqrt{bB^{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}$
$\quad+2\min\left\{\sqrt{\frac{b(D+1)^{2}}{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}},\left(bB^{4}C_{x}^{2}C_{y}^{2}(D+1)^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}\right\}-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})},$ (13)
where we used the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$. Therefore, substituting (12) and (E.3) into (11) gives
$\mathcal{R}_{T}$ $\leq 4BC_{x}C_{y}D+\frac{\sqrt{2}}{3}\log\frac{1}{\delta}+\left(\sqrt{2(1-a)\log\frac{1}{\delta}}+2\sqrt{bB^{2}}\right)\sqrt{\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}$
$+\min\left\{\sqrt{2b(D+1)^{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}},2\left(bB^{4}C_{x}^{2}C_{y}^{2}(D+1)^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}\right\}-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}$
$\leq 4BC_{x}C_{y}D+\frac{\sqrt{2}}{3}\log\frac{1}{\delta}+\frac{\left(\sqrt{(1-a)\log\frac{1}{\delta}}+\sqrt{2bB^{2}}\right)^{2}}{a}$
$+\min\left\{\frac{b(D+1)^{2}}{a},\frac{3}{2}\left(a^{-1}bB^{4}C_{x}^{2}C_{y}^{2}(D+1)^{2}T\right)^{1/3}\right\},$
@@
Appendix F Details omitted from Section 5
@@ F.1 Common analysis
We provide the analysis that is commonly used in the proofs of Theorems 5.1 and 5.2. Although we use $\widehat{\bm{G}}_{t}$ as a notation for the estimator for convenience in this subsection, the same argument applies equally to $\widetilde{\bm{G}}_{t}$. We use ODAFTRL with the AdaHedgeD update in Section E.1 as $\mathsf{ALG}$. Here, we recall that $\mathbb{E}[\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}]\leq\mathbb{E}[\sum_{t=1}^{T}{\langle\widehat{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}]$ from the convexity of $S_{t}$ and the unbiasness of $\widehat{\bm{G}}_{t}$. From Lemma E.1, it holds that
$\sum_{t=1}^{T}{\langle\widehat{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\leq 2\left(2\max_{t\in[T]}a_{t-D:t-1}+\sqrt{\sum_{t=1}^{T}{a_{t}^{2}}+B^{2}b_{t}}\right),$ (14)
@@
$a_{t}=B\min\{\|\widehat{\bm{G}}_{t-D:t}\|_{\mathrm{F}},\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}\}\quad\mbox{and}\quad b_{t}\leq\min\left\{\frac{1}{2}\|\widehat{\bm{G}}_{t-D:t}\|_{\mathrm{F}}^{2},\|\widehat{\bm{G}}_{t-D:t}\|_{\mathrm{F}}\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}\right\}.$
@@
$\mathbb{E}\left[\max_{t\in[T]}a_{t-D:t-1}\right]\leq B\mathbb{E}\left[\max_{t\in[T]}\sum_{s=t-D}^{t-1}\|\widehat{\bm{G}}_{s}\|_{\mathrm{F}}\right],$ (15)
and thus
$\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widehat{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]$ $\leq 2\left(2\mathbb{E}\left[\max_{t\in[T]}a_{t-D:t-1}\right]+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{a_{t}^{2}}\right]}+B\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]}\right)$
$\!\leq 2B\left(2\mathbb{E}\left[\max_{t\in[T]}\sum_{s=t-D}^{t-1}\|\widehat{\bm{G}}_{s}\|_{\mathrm{F}}\right]+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]}\right),$ (16)
@@
$\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t-D:t}\|_{\mathrm{F}}\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}}\right]\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}\sum_{s=t-D}^{t-1}\|\widehat{\bm{G}}_{s}\|_{\mathrm{F}}}\right]$
$=\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\mathbb{E}\left[\sum_{t=1}^{T}{\mathbb{E}_{t}\left[\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}\right]\sum_{s=t-D}^{t-1}\|\widehat{\bm{G}}_{s}\|_{\mathrm{F}}}\right],$ (17)
@@
$\mathbb{E}\left[\mathcal{R}_{T}\right]\leq 4BC_{x}C_{y}D+\left(\frac{4bB}{a}+1\right)\sqrt{KT}+2BC_{x}C_{y}\sqrt{DT}=O(\sqrt{(K+D)T}).$
@@
$\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\mathbb{E}\left[\sum_{t=1}^{T}{\mathbb{E}_{t}\left[\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}\right]\sum_{s=t-D}^{t-1}\|\widehat{\bm{G}}_{s}\|_{\mathrm{F}}}\right]$
$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+C_{x}C_{y}\mathbb{E}\left[\sum_{t=1}^{T}{\sum_{s=t-D}^{t-1}\mathbb{E}_{s}\left[\|\widehat{\bm{G}}_{s}\|_{\mathrm{F}}\right]}\right]$
$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+C_{x}^{2}C_{y}^{2}DT,$ (18)
where the second and third inequality follow from the inequality $\mathbb{E}_{t}\left[{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}}\right]=\|\bm{G}_{t}\|_{\mathrm{F}}\leq C_{x}C_{y}$. Hence, from $\mathbb{E}_{t}\left[{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}}\right]\leq C_{x}C_{y}$, (F.1) and (F.2), it holds that
$\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widehat{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]$ $\leq 2B\left(2C_{x}C_{y}D+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+C_{x}^{2}C_{y}^{2}DT}}\right)$
$\leq 4BC_{x}C_{y}D+4B\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+2BC_{x}C_{y}\sqrt{DT}$
$\leq 4BC_{x}C_{y}D+4B\sqrt{\frac{bK}{q}\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}+2BC_{x}C_{y}\sqrt{DT},$ (19)
@@
$\mathbb{E}_{t}\left[{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]=\frac{\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}{p_{t}(\bm{y}_{t})}\leq\frac{K}{q}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq\frac{bK}{q}S_{t}(\bm{W}_{t}).$
@@
$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widehat{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
$\leq 4BC_{x}C_{y}D+4B\sqrt{\frac{bK}{q}\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}+2BC_{x}C_{y}\sqrt{DT}-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
$\leq 4BC_{x}C_{y}D+\frac{4bB^{2}}{a}\frac{bK}{q}+2BC_{x}C_{y}\sqrt{DT}+qT,$
where the first inequality follows from Assumption 3.2, the second inequality follows from the convexity of $S_{t}$ and the unbiasedness of $\widehat{\bm{G}}_{t}$, and the last inequality follows from $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$. Finally, from $q=B\sqrt{K/T}$, we obtain
$\mathbb{E}\left[\mathcal{R}_{T}\right]\leq 4BC_{x}C_{y}D+\left(\frac{4bB}{a}+1\right)\sqrt{KT}+2BC_{x}C_{y}\sqrt{DT},$
@@
$\mathbb{E}[\mathcal{R}_{T}]=4BC_{x}C_{y}(2D+\sqrt{DT})+\frac{8bB^{2}}{a}+O\left(\omega^{1/3}D^{1/3}T^{2/3}\right).$
@@
$\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}}\right]$ $=\mathbb{E}_{t}\left[{\|(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\widetilde{\bm{y}}_{t})\bm{x}_{t}^{\top}\|_{\mathrm{F}}}\right]$
$\leq\mathbb{E}_{t}\left[{\|\bm{G}_{t}\|_{\mathrm{F}}+C_{x}\|\bm{y}_{t}-\widetilde{\bm{y}}_{t}\|_{2}}\right]\leq\|\bm{G}_{t}\|_{\mathrm{F}}+C_{x}\mathbb{E}_{t}\left[{\|\bm{y}_{t}\|_{2}+\|\widetilde{\bm{y}}_{t}\|_{2}}\right]$
@@
$\leq\|\bm{G}_{t}\|_{\mathrm{F}}+C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q}\leq 2C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q},$ (20)
where the first inequality follows from $C_{x}\geq\|\bm{x}_{t}\|_{2}$, the third inequality follows from $C_{y}\geq\|\bm{y}_{t}\|_{2}$, the fourth inequality follows from Jensen’s inequality, and the fifth inequality follows from Lemma D.6. Thus, combining (F.1) with the last inequality, we have
@@
$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\left(2C_{x}C_{y}+\sqrt{\frac{C_{x}^{2}\omega}{q}}\right)\mathbb{E}\left[\sum_{t=1}^{T}{\sum_{s=t-D}^{t-1}\mathbb{E}_{s}\left[\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}\right]}\right]$
$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+DT\left(2C_{x}C_{y}+\sqrt{\frac{C_{x}^{2}\omega}{q}}\right)^{2}.$ (21)
Hence, from (F.1), (F.3),and (21), we have
@@
$\leq 2B\left(2\mathbb{E}\left[\max_{t\in[T]}\sum_{s=t-D}^{t-1}\|\widehat{\bm{G}}_{s}\|_{\mathrm{F}}\right]+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]}\right)$
$\leq 2B\left(4C_{x}C_{y}D+2C_{x}D\sqrt{\omega/q}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]}\right)$
$\leq 8BC_{x}C_{y}D+4BC_{x}D\sqrt{\omega/q}+4B\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+2B\left(2C_{x}C_{y}+C_{x}\sqrt{\omega/q}\right)\sqrt{DT}$
$\leq 8BC_{x}C_{y}D+4BC_{x}D\sqrt{\omega/q}$
$\qquad+4B\sqrt{2\sum_{t=1}^{T}{\left(bS_{t}(\bm{W}_{t})+\frac{C_{x}^{2}\omega}{q}\right)}}+2B\left(2C_{x}C_{y}+C_{x}\sqrt{\omega/q}\right)\sqrt{DT},$ (22)
where the first inequality follows from (F.1), the second inequality follows from (F.3), the third inequality follows from (21) and the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$, and the last inequality follows from Lemma D.7. Therefore, combining all the above arguments yields
@@
$\leq 8BC_{x}C_{y}D+4B\left(\sqrt{2b\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}+C_{x}\sqrt{2\omega T/q}\right)+4BC_{x}D\sqrt{\omega/q}$
$\qquad+2B\left(2C_{x}C_{y}+C_{x}\sqrt{\omega/q}\right)\sqrt{DT}-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
$\leq 4BC_{x}C_{y}(2D+\sqrt{DT})+\frac{8bB^{2}}{a}+4BC_{x}D\sqrt{\omega/q}+2BC_{x}\big{(}\sqrt{D}+2\sqrt{2}\big{)}\sqrt{\omega T/q}+qT,$